OpenStudy (anonymous):

Solve 4x2 = 12x + 40.

OpenStudy (callisto):

4x^2 = 12x + 40 x^2 = 3x +10 x^2 -3x - 10 =0 (x-5)(x+2) =0 (x-5)=0 or (x+2) =0 Can you solve it?

OpenStudy (anonymous):

no :( what do i do after this?

OpenStudy (callisto):

Solve the 2 equations ._. First, solve x-5 =0 .... Then, x+2 = 0 ....